The Specific Heat of Aluminium is 0.89 J/gC. How much heat must be applied to a 10 gram sample to raise its temperature 2 degrees?
Answer #1

q=m*cp*DeltaT

q is the amount of heat required

m is the mass of the substance

cp is the specific heat of the substance

DeltaT is the temperature rise or decrease

q=10*0.89*2=17.8J